OpenStudy (azureflame241):

I need help solving: (x/(x+4))+(1/x)-4=16/((x^2)+4x)

OpenStudy (azureflame241):

Heres a more friendly view of the equation: \[\frac{ x }{ x+4 }+\frac{ 1 }{ x }-4=\frac{ 16 }{ x ^{2}+4x }\]